find the slope and the y-intercept of the line. Write fractional answers in lowest terms. 2y+7x+7=19+7x
Answer by nerdybill(7384)

2y+7x+7=19+7x
subtractiong 7x from both sides:
2y+7=19
subtracting 7 from both sides:
2y=12
dividing both sides by 2
y = 6
.
slope is 0
y-intercept is at (0,6)